# Chaotic interaction between dark matter and dark energy

### v5

@@ Abstract
In this study, we consider dark matter and dark energy as grand-canonical systems which are open, non-equilibrium coupled, and interacting systems. For the first time, we propose a new more realistic interaction scheme to explain dynamics between coupled interacting thermodynamic systems. Based on this new interaction schema, we propose new theorems to define the interactions including mutual and self-interactions. We proved the theorems based on the energy conservation law of thermodynamics. Furthermore, we obtain new coupled equations using the theorems. We numerically solved the interaction equations and obtained phase space diagrams and Lyapunov exponents. We show that the interaction between dark matter and dark energy is chaotic. We conclude that these theorems and results can be generalized to all coupled interacting particle and thermodynamic systems. Finally, we evaluated that the obtained results indicate the existence of a new law in physics. Additionally, based on the results, we gave physical definitions of basic concepts such as chaos and self-organization and pointed out the existence of hidden symmetries and variables in interacting systems.
@@ I Introduction
Although the physical origin of dark matter and dark energy has not been explained yet, it is known that they play an important role in the dynamics of the universe. For this reason, understanding the interaction of all the components that make up the universe and the results arising from these interactions is one of the most important problems of cosmology. To tackle this problem, many interacting models based on dark matter and dark energy interaction have been used to solve the singularity and cosmic coincidence problems [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 12, 13, 30, 31, 32, 33, 34, 35, 36, 37].
@@
In 2018, we proposed Chaotic Universe Theory based on interactions between the components such as matter, dark matter, and dark energy [38]. In this study, we showed that the interaction dynamics of these components are chaotic. We suggested that the universe is evolving chaotically within the cosmic time-line. We concluded that this new cosmological model has the potential to solve many complications of the big-bang cosmology such as singularities, cosmic coincidence, emergence of galaxies, fractal formation of the universe from micro to macro scale, and the future and the end of the universe. Additionally, in another study, we showed that dark energy and dark matter interaction introduce a novel mechanism to explain late time inflation i.e., the transition from matter-dominated era to dark energy-dominated era of the universe [39]. The first time, we obtained a hybrid scale factor containing both eras and showed that this transition appears around 9.8-10 Gyears in the cosmic time-line. Furthermore, we found that the parameters of the new hybrid scale factor are precisely consistent with the theoretical and observational results for the matter and dark energy-dominated eras. Finally, the importance of the dark matter and dark energy interactions has been thoroughly discussed in Ref.  [40].
@@
The results presented above emphasize that developing interaction, particularly, chaotic interaction models may have great importance in cosmology. In Ref. [38], we obtained the chaotic interactions between matter, dark matter, and dark energy inspired by Lotka-Volterra-type interactions. However, even though the Lotka-Volterra-type interaction equations of dark matter and dark energy models can cause chaotic behavior under some special conditions [38, 41], these equations may not fully describe the interaction between dark matter and dark energy. Therefore, in this study, we will propose a new interaction schema for dark matter and dark energy interactions, supporting the main claim of Ref.  [38]. Furthermore, we will consider dark matter and dark energy as an open and interacting thermodynamic system in a new interaction schema, unlike standard statistical thermodynamics. Based on this new theoretical framework, we will show that interaction equations can be derived from the first law of thermodynamics. To obtain a full description of the interaction we will suggest a new complementary equation inspired by graph theory. Finally, based on these theorems, we will carry out new interaction equations that describe the interaction between dark matter and dark energy. We show that the interaction between dark matter and dark energy is chaotic. Our results might give a critical contribution not only to cosmology but also to statistical thermodynamics and the understanding of nature and living systems.
@@ III.1 Mutual Interactions
By using Lemma 1, the internal energy change in Eq.9 can be written as
@@ III.2 Self-Interactions
Figure 3: The vectorial attractor fields and torque behavior on the self-interacting loops. $\Gamma_{1,2}$ denotes non-holonomic variables in Eq.6. $\vec{A}_{1,2}$ are the attractor vectorial fields caused by driven force $\vec{F}_{1,2}$ and $\vec{\tau}_{1,2}$ correspond to torque on the loops.
@@
where loop radius $r_{1,2}$ characterize the velocity magnitude of the densities and $A_{1,2}$ are the vectorial fields on the loops. The integrals in Eq.16 can be written as
@@
where $M_{1,2}$ are the positive constants. As a result, combining Eqs.15, 16 and 17, we obtain
@@ IV Numerical Results
To analyze the dynamical behavior of Eq.III.3, we solved these equations numerically by writing FORTRAN 90 code based on the Runga-Kutta method and the linearized algorithm [46]. Phase trajectories are given in Figs.4-6.
@@
We compute the Lyapunov exponent by using the Wolf algorithm and give the numerical results in Fig.7. As seen from Fig.7, there are three Lyapunov exponents $\lambda_{i}$, $i=1,2,3$ for the control parameter $q$. Positive Lyapunov exponents indicate the chaotic behavior of the system. One can see that two exponents (red and green curves) have positive values depending on the control parameter $q$.
@@
Lyapunov spectrum in Fig.7 shows that the system in Eq.III.3 behaves chaotic. Based on these numerical results we conclude that the interaction between dark matter and dark energy is chaotic.
@@ V Discussion and Some Concluding Remarks
We propose a new interaction scheme that represents the interaction of dark matter and dark energy. We have demonstrated that interactions should be defined in two ways. The first is the mutual interaction between the two systems and the second is the self-interaction that occurs on each system. These two interaction pictures are beyond what has been known so far and offer a new approach to the interaction mechanisms of interacting systems. Based on the new interaction schema, we introduce several new concepts such as energy transfer velocity, vectorial attractor fields, vectorial attractor torques which are new concepts for the non-equilibrium statistical mechanics and thermodynamics. Furthermore, we introduce new theorems to define the mutual and self-interactions.
@@
It is very important to mention another crucial point. The interaction scheme in Fig.3, Theorem 1, Theorem 2 and interaction equations in Eq. III.3 and all the results we obtained here are valid not only for cosmology but also for all coupled non-equilibrium thermodynamic systems. The results also show that the dynamics of all coupled, interacting, and transforming thermodynamic systems are chaotic both far from and close to the equilibrium. This is also another significant result of the present study. Additionally, these results may indicate the presence of new thermodynamics laws.
@@ VII Further Studies
One can be seen from Section VI, this work contains many open problems. For now, we have examined our proposed interaction scheme for a few special cases of open and closed thermodynamic systems [47]. Furthermore, we are now continuing our detailed work on some of these open problems. Particularly, following the concept of this work, it is of great importance to continue studies on the new interaction scheme and its effects on cosmology. Approaching the problem from this perspective, we will discuss basic cosmology problems such as singularity, flatness, monopole, transitions between different epochs, entropy, complexity, and the future of the universe by resolving the interaction equations Eq. (III.3) using standard cosmology $\Lambda$CDM data. In this way, we may have the chance to test the consistency of the presented model. On the other hand, testing the ideas proposed for particle and thermodynamic systems is also crucial. This is another important open problem.
@@ Acknowledgements
I am grateful to the anonymous referee for the many valuable constructive corrections, suggestions, and contributions that played an important role in improving the manuscript. This work was supported by İstanbul University. Research Project: FYO-2021-38105 which is entitled “Investigation of the Cosmic Evolution of the Universe”.

### v6

@@ Abstract
In this study, we consider dark matter and dark energy as grand-canonical systems which are open, non-equilibrium, coupled, and interacting systems. For the first time, we propose a new more realistic interaction scheme to explain dynamics between coupled interacting thermodynamic systems. Based on this new interaction schema, we propose new theorems to define the interactions including mutual and self-interactions. We proved the theorems based on the energy conservation law of thermodynamics. Furthermore, we obtain new coupled equations using the theorems. We numerically solved the interaction equations and obtained phase space diagrams and Lyapunov exponents. We show that the interaction between dark matter and dark energy is chaotic. We conclude that these theorems and results can be generalized to all coupled interacting particle and thermodynamic systems. Finally, we evaluated that the obtained results indicate the existence of a new law in physics. Additionally, based on the results, we gave physical definitions of basic concepts such as chaos and self-organization and pointed out the existence of hidden symmetries and variables in interacting systems.
@@ I Introduction
The physical origins of dark matter and dark energy are not yet understood. This is among the important problems of the particle physics and cosmology. However, we know that these components interact at least through gravity. Therefore, the interactions between these components may play an important role in the evolution of the universe. Recently, there has been an increasing number of experimental, observational and theoretical studies aimed at understanding the physical origin of dark matter and dark energy and determining the interactions between them. For instance, there are many publications in the literature that model the interactions in the dark sector to solve singularity, cosmic coincidence, Hubble tension problems and etc [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 12, 13, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50].
@@
In addition to the above studies, in 2018, we proposed Chaotic Universe Theory based on interactions between the components such as matter, dark matter, and dark energy [51]. In this study, we showed that the interaction dynamics of these components are chaotic. We suggested that the universe is evolving chaotically within the cosmic time-line. We concluded that this new cosmological model has the potential to solve many complications of the big-bang cosmology such as singularities, cosmic coincidence, emergence of galaxies, fractal formation of the universe from micro to macro scale, and the future and the end of the universe. Additionally, in another study, we showed that dark energy and dark matter interaction introduce a novel mechanism to explain late time inflation i.e., the transition from matter-dominated era to dark energy-dominated era of the universe [52]. The first time, we obtained a hybrid scale factor containing both eras and showed that this transition appears around 9.8-10 Gyears in the cosmic time-line. Furthermore, we found that the parameters of the new hybrid scale factor are precisely consistent with the theoretical and observational results for the matter and dark energy-dominated eras. Finally, the importance of the dark matter and dark energy interactions has been thoroughly discussed in Ref.[53].
As we have briefly summarized above, interaction models seem to have significant potential to solve some important problems in cosmology. This is our main motivation in this study. On the other hand, in addition to this main motivation, we have several more important motivations that are the focus of this study. First of all, the nature we live in, the universe, is non-linear. This non-linearity leads to chaotic dynamics. Second, our universe has a fractal geometry from micro to macro scale. It is expected that the dynamics that lead to this geometric elegance are chaotic. This is the intuitive reason why we expect the interactions of the components in the universe to be chaotic. Third, in Ref. [51], we obtained the chaotic interactions between matter, dark matter, and dark energy inspired by Lotka-Volterra-type interactions. However, even though the Lotka-Volterra-type interaction equations of dark matter and dark energy models can cause chaotic behavior under some special conditions [51, 54], these equations may not fully describe the interaction between dark matter and dark energy. More importantly, we need newer approaches to the physics of interacting systems that go beyond standard approaches and reveal the deeper physics and enable understanding of nature. In this study, we will try to answer these motivations.
Therefore, here, we will propose a new interaction schema for dark matter and dark energy interactions, supporting the main claim of Ref.[51] and to answer to the above motivations. To be able to do this, firstly, we will consider dark matter and dark energy as an open and interacting thermodynamic system in a new interaction schema, unlike standard statistical thermodynamics. We have proposed two theorems as the theoretical counterpart of this interaction scheme. One of them is mutual interaction and the other is self-interaction.
@@
Based on this new theoretical framework, we will show that interaction equations can be derived from the first law of thermodynamics. To obtain a full description of the interaction we will suggest a new complementary equation inspired by graph theory. Finally, based on these theorems, we will carry out new interaction equations that describe the interaction between dark matter and dark energy. We will show that the interaction between dark matter and dark energy is chaotic. On the other hand, we aim to propose a new law of physics, inspired by the possibility that the obtained results may be universal for all thermodynamic systems. Additionally, we will introduce and explore several new conceptual definitions. We believe that our results might give a critical contribution not only to cosmology but also to statistical thermodynamics and the understanding of nature and living systems.
@@ III.1 Mutual Interactions
By using Lemma 1, the internal energy change in Eq.(9) can be written as
@@ III.2 Self-Interactions
Figure 3: The vectorial attractor fields and torque behavior on the self-interacting loops. $\Gamma_{1,2}$ denotes non-holonomic variables in Eq.(6). $\vec{A}_{1,2}$ are the attractor vectorial fields caused by driven force $\vec{F}_{1,2}$ and $\vec{\tau}_{1,2}$ correspond to torque on the loops.
@@
where loop radius $r_{1,2}$ characterize the velocity magnitude of the densities and $A_{1,2}$ are the vectorial fields on the loops. The integrals in Eq.(16) can be written as
@@
where $M_{1,2}$ are the positive constants. As a result, combining Eqs.(15), (16) and (17), we obtain
@@ IV Numerical Results
To analyze the dynamical behavior of Eq.(III.3), we solved these equations numerically by writing FORTRAN 90 code based on the Runga-Kutta method and the linearized algorithm [59]. Phase trajectories are given in Figs.4-6.
@@
Figure 7: Chaotic attractors in $x_{1}-x_{2}-x_{3}$ phase plane for the control parameter $q=3.46$.
This situation can be observed more clearly in Fig.,7. In this phase space, $x_{1}$, $x_{2}$, and $x_{3}$ are plotted against each other for the same initial conditions and $q$ value. From Fig.,7, it is evident that the trajectories within the attractors do not intersect. If the trajectories were to intersect, the solutions would no longer be expected to exhibit chaotic behavior. The crucial point here is the demonstration of the existence of chaotic attractors. It is possible to reproduce such plots for various $q$ values. Whether these attractors are indeed chaotic can be further clarified using the Lyapunov exponent.
Additionally, we compute the Lyapunov exponent by using the Wolf algorithm and give the numerical results in Fig.8. As seen from Fig.8, there are three Lyapunov exponents $\lambda_{i}$, $i=1,2,3$ for the control parameter $q$. Positive Lyapunov exponents indicate the chaotic behavior of the system. One can see that two exponents (red and green curves) have positive values depending on the control parameter $q$.
@@
Lyapunov spectrum in Fig.8 shows that the system in Eq.(III.3) behaves chaotic. Based on these numerical results we conclude that the interaction between dark matter and dark energy is chaotic.
It can be observed from Fig.8 that the Lyapunov exponent has a positive value for $q<25$, while it fluctuates around zero for $q>25$. A more detailed analysis may be required to examine the specifics; however, this is not essential for the current discussion. It is sufficient to note that a positive Lyapunov exponent indicates chaotic behavior in the system. Nevertheless, the system may not exhibit chaos for all values of $q$. The parameter $q$ represents the difference between $\Gamma_{1}$ and $\Gamma_{2}$. As this difference increases beyond $q>25$, the system may transition out of chaotic behavior. Specifically, scenarios where $\Gamma_{1}$ $\gg$ $\Gamma_{2}$ or $\Gamma_{1}$ $\ll$ $\Gamma_{2}$ could lead the nonlinear system to produce linear solutions. Consequently, the interaction between dark matter and dark energy would no longer exhibit chaotic dynamics. To analyze this phenomenon from a cosmological perspective, it is crucial to determine the appropriate range of $q$. Achieving this requires precise fitting of parameters such as the dark matter and dark energy densities, the equation of state parameter $\omega$, cosmological quantities like the scale factor $a(t)$, and the known data from the $\Lambda$CDM model. This topic will be revisited in Section VII.
@@ V Discussion and Some Concluding Remarks
We propose a new interaction scheme that represents the interaction of dark matter and dark energy. We have demonstrated that interactions should be defined in two ways. The first is the mutual interaction between the two systems and the second is the self-interaction that occurs on each system. These two interaction pictures are beyond what has been known so far and offer a new approach to the interaction mechanisms of interacting systems. Based on the new interaction schema, we introduced several new concepts such as energy transfer velocity, vectorial attractor fields, vectorial attractor torques which are new concepts for the non-equilibrium statistical mechanics and thermodynamics. Furthermore, we introduced new theorems to define the mutual and self-interactions.
@@
It is very important to mention another crucial point. The interaction scheme in Fig.3, Theorem 1, Theorem 2 and interaction equations in Eq. (III.3) and all the results we obtained here are valid not only for cosmology but also for all coupled non-equilibrium thermodynamic systems. The results also show that the dynamics of all coupled, interacting, and transforming thermodynamic systems are chaotic both far from and close to the equilibrium. This is also another significant result of the present study. Additionally, these results may indicate the presence of new thermodynamics laws.
@@ VII Further Studies
In this work, we propose a new interaction scheme and new definitions for thermodynamic systems. Using the scheme, we show that the interactions between dark matter and dark energy must be chaotic. Furthermore, the proposed new interaction scheme offers a new perspective on physics. Therefore, the model may have applications in many areas. On the other hand, one can see that the present study contains many open problems. For instance, we recently examined the interactions between two thermodynamic systems under specific conditions using this scheme [60].
In particular, one important open problem is to understand whether the model is compatible with standard gravity theories. Therefore, we are currently resolving this interaction scheme by considering the well-known results of $\Lambda$CDM. One of our most important goals is in Eq. (III.3) to see whether our proposed interaction model will work in accordance with the known results of cosmology. Currently, we are trying to determine the working range of the interaction model by fitting the model to the well-known results of $\Lambda$CDM. Furthermore, we are discussing the scale factor, dark matter and dark energy densities, state parameters and other known cosmological quantities using the model. Similarly, some fundamental problems of cosmology such as singularity, flatness, monopole, transitions between different epochs, entropy, complexity, and the future of the universe are among our immediate goals.
Additionally, discussing the proposed model within the framework of thermodynamics and particle physics and understanding its application areas may provide interesting results. On the other hand, experimental verification of this proposed interaction scheme is of great importance. Studying these open problems may contribute to the development of new insights and is of great importance in terms of the validation of the model.
@@ Acknowledgements
I am grateful to the anonymous referee of IJTP for valuable comments and suggestions that improved the quality of the work. I am grateful to the anonymous referee in the previous submission for the many valuable constructive corrections, suggestions, and contributions that played an important role in improving the manuscript. Finally, I thanks İstanbul University since this work was supported by İstanbul University. Research Project: FYO-2021-38105 which is entitled “Investigation of the Cosmic Evolution of the Universe”.